OpenStudy (anonymous):

64^(x)=16^(x-1) x=-2 x=-1 x=-1/4 x=-1/3

sam (.sam.):

64^(x)=16^(x-1) 4^(3(x))=4^(2(x-1)) 3(x)=2(x-1) 3(x)=2x-2 3x=2x-2 x=-2